OpenStudy (anonymous):

You graph a system of equations and the two lines overlap, or lie right on top of one another. Explain what this tells you about the: solution to the system the two equations the result if the system was solved algebraically @Compassionate

OpenStudy (anonymous):

There's no solution if the lines overlap, right? Or is that just my tired brain talking.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you must be tired because two lines that overlap will intersect an infinite amount of times

jimthompson5910 (jim_thompson5910):

so there are infinitely many solutions (all on the lines that overlap) this system is then known to be a dependent system (since one equation depends on the other)

OpenStudy (anonymous):

Oh jeez. Way off. What about #2?

jimthompson5910 (jim_thompson5910):

well the two equations would be very similar since one equation would just be a scalar multiple of the other

jimthompson5910 (jim_thompson5910):

ex: x+y = 10 and 2x+2y = 20 form a dependent system of equations

jimthompson5910 (jim_thompson5910):

since you're doubling everything in x+y=10 to get 2x+2y=20

OpenStudy (anonymous):

and the third?

jimthompson5910 (jim_thompson5910):

at the very end, you would get 0=0, which is always true that leads to the fact that there are infinitely many solutions

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw